OpenStudy (atomicreaper):

In the equation -3(y-2)^2/3+29=-19 the values of y are?

OpenStudy (jango_in_dtown):

put y-2=z^3. then we have -3z^2+29=-19 this implies -3z^2=-48 this imply z^2=16 this imply z=4, -4 this imply y-2=64,-64 this imply y=66,-62

OpenStudy (jango_in_dtown):

@AtomicReaper

OpenStudy (jango_in_dtown):

put y-2=z^3. then we have -3z^2+29=-19 this implies -3z^2=-48 this imply z^2=16 this imply z=4, -4 this imply y-2=64,-64 this imply y=66,-62